# Aligning Cyber Space with Physical World: A Comprehensive Survey on Embodied AI

### v7

@@ V Embodied Interaction
Embodied interaction tasks refer to scenarios where agents interact with humans and environment in a physical or simulated space. The typical embodied interaction tasks are Embodied Question Answering (EQA) and embodied grasping.
@@ V-A1 Datasets
Conducting robot experiments in real environments is often constrained by scenarios and robot hardware. As virtual experimental platforms, simulators offer suitable environmental conditions for constructing embodied question answering datasets. Training and testing models on datasets created in simulators significantly reduce experimental costs and enhance the success rate of deploying models on real machines. We briefly introduce several embodied question answering datasets, which are summarized in Table IX.
TABLE IX: Comparison of Different EQA Datasets.
@@
EQA v1[259] is the first dataset designed for EQA. Built on synthetic 3D indoor scenes from the SUNCG dataset[95] within the House3D[269] simulator, EQA v1 comprises four types of questions: location, color, color_room, and preposition. It features over 5,000 questions distributed across more than 750 environments. The questions are constructed via functional program execution, using templates to select and combine basic operations.
Similar to EQA v1, MT-EQA[260] is built in House3D using SUNCG by executing functional programs consisting of some basic operations. However, it further extends the single-object question answering task to a multi-object setting. Six types of questions are designed, involving the comparison of color, distance, and size between multiple objects. The dataset contains 19,287 questions in 588 environments.
MP3D-EQA[261] is built on a simulator developed based on MINOS[270] using the Matterport3D dataset[271], expanding the question-answering task to a realistic 3D environment. Referring to EQA v1, MP3D-EQA utilizes three types of templates: location, color, and color_room, generating a total of 1,136 questions in 83 home environments.
IQUAD V1[262] is built upon AI2-THOR and consists of three types of questions: existence, counting, and spatial relationships. It uses a set of templates written down a priori to generate more than 75,000 multiple choice questions, each accompanied by a unique scene configuration. Unlike other datasets, answering IQUAD V1 questions requires the agent to have a good understanding of affordances and interact with the dynamic environment.
VideoNavQA[263] decouples the visual reasoning from the navigation aspect of the EQA problem. In this task, the agent accesses videos corresponding to exploration trajectories with sufficient information to answer questions. Still referring to EQA v1, VideoNavQA generates questions according to functional, template-style representation. It also renders shortest trajectories to simulate near-optimal navigation paths, creating videos corresponding to what an agent would see while exploring the environment. VideoNavQA generates about 101,000 pairs of videos and questions in the House3D environment using SUNCG, covering 28 types of questions belonging to 8 categories such as existence, counting, and localization.
SQA3D[264] simplifies protocol (QA only) while still preserving the function of benchmarking embodied scene understanding, enabling more complex, knowledge-intensive questions and a much larger scale of data collection. Specifically, SQA3D offers a dataset with about 6,800 unique situations, 20,400 descriptions, and 33,400 diverse reasoning questions for these situations based on ScanNet[272] scenes.
Unlike previous datasets that explicitly specify target objects in questions, K-EQA[265] features complex questions with logical clauses and knowledge-related phrases, requiring prior knowledge to answer. It is built in AI2Thor and includes four types of questions: existence, counting, enumeration, and comparison. Each entity is mapped to a knowledge base and a knowledge graph is further constructed. In this work, the templates provided in IQA[262] and MT-EQA are extended to a set of grammars. After specifying objects and logical relationships, knowledge graphs, scene graphs, etc. are introduced to generate questions and compute the ground truth answer. The resulting K-EQA dataset consists of 60,000 questions across 6000 different environment setups.
OpenEQA[266] is the first open-vocabulary dataset for EQA, supporting both episodic memory and active exploration cases. The episodic memory EQA (EM-EQA) tasks involve an agent developing an understanding of the environment from its episodic memory to answer questions, similar to VideoNavQA. In active EQA (A-EQA) tasks, the agent answers questions by taking exploratory actions to gather necessary information. Using ScanNet and HM3D[273], human annotators constructed over 1,600 high-quality questions from more than 180 real world environments in Habitat.
Utilizing GPT4-V, HM-EQA[267] is constructed in the Habitat simulator using HM3D. It includes 500 questions across 267 different scenes, which can be roughly categorized into identification, counting, existence, status, and location. For consistency, each question has four multiple choices.
S-EQA[268] leverages GPT-4 in VirtualHome for data generation and employs cosine similarity calculations to decide whether to retain the generated data, thereby enhancing dataset diversity. In S-EQA, answering questions requires the assessment of a collection of consensus objects and states to reach an existential “Yes/No” answer.
@@ V-A2 Methods
The EQA task was first proposed by Das et al.[259]. In their work, the agent consists of four main modules: vision, language, navigation, and answering. These modules are primarily constructed using traditional neural building blocks: Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs). They undergo training in two phases. Initially, the navigation and response modules are trained independently on automatically generated expert navigation demonstrations using imitation or supervised learning. Subsequently, in the second phase, the navigation architecture is fine-tuned using policy gradients. Some subsequent works[274, 275] retained modules like the question answering module proposed by Das et al.[259] and improved the model. Additionally, Wu et al.[275] proposed integrating the navigation and QA modules into a unified SGD training pipeline for joint training, thereby avoiding employing deep reinforcement learning to simultaneously train the separately trained navigation and question answering modules.
There are also some works that attempted to increase the complexity and completeness of question answering tasks. From the perspective of task singularity, several works[260, 276] expanded the task to include multiple objectives and multi-agent, respectively, making it necessary for the model to store and integrate the information obtained by the agent’s exploration through methods such as feature extraction and scene reconstruction. Taking into account the interaction between the agent and the dynamic environment, Gordon et al.[262] introduced the Hierarchical Interactive Memory Network. Control alternates between the planner, responsible for task selection, and the low level controllers, which carry out task execution. During this process, an Egocentric Spatial GRU (esGRU) is utilized to store spatial memory, enabling the agent to navigate and provide answers. There is also a limitation in previous works where agents are unable to use external knowledge to answer complex questions and lack knowledge of the explored parts of the scene. To address this, Tan et al.[265] proposed a framework that leverages the neural program synthesis method and the table converted from the knowledge and 3D scene graphs, allowing the action planner to access object-related information. Additionally, an approach based on Monte Carlo Tree Search (MCTS) is used to determine the next location for the agent to move to.
LLMs/VLMs Methods. In recent years, LLMs and VLMs have made continuous progress and demonstrated outstanding capabilities across various fields. Consequently, researchers attempt to apply these models to solve embodied question-answering tasks without any additional fine-tuning.
@@
Majumdar et al.[266] explored using LLMs and VLMs for episodic memory EQA (EM-EQA) task and Active EQA (A-EQA) task. For EM-EQA task, they considered Blind LLMs, Socratic LLMs with language descriptions of the episodic memory, Socratic LLMs with descriptions of the constructed scene graph, and VLMs processing multiple scene frames. The A-EQA task extended EM-EQA methods with frontier-based exploration (FBE)[277] for problem-independent environment exploration. Some other works[267, 278] also employed frontier-based exploration method to identify areas for subsequent exploration and to build semantic maps. They ended the exploration early utilizing conformal prediction or image-text matching to avoid over-exploration. Patel et al.[279] emphasized the question answering aspect of the task. They leveraged multiple LLM-based agents to explore the environment and enable them to independently answer questions with “yes” or “no” answers. These individual responses are utilized to train a Central Answer Model, responsible for aggregating the responses and generating robust answers.
V-A3 Metrics
The performance is usually assessed based on two aspects: navigation and question answering. In the navigation, many works adhered to the approach introduced by Das et al.[259] and utilized indicators like the distance to the target object upon completion of navigation ($d_{T}$), the change in distance to target from initial to final position ($d_{\varDelta}$) and the smallest distance to the target at any point in the episode ($d_{min}$) to evaluate the performance of the model. They are tested at 10, 30, or 50 actions away from the target. There are also works that measured it based on indicators such as trajectory length, intersection-over-union score for target object ($IoU$), etc. For question answering, the evaluation mainly involves mean rank ($MR$) of the ground-truth answer in the answer list and accuracy of the answers. Recently, Majumdar et al.[266] introduced the concept of an aggregate LLM-based correctness metric (LLM-Match) to evaluate the accuracy of open-vocabulary answers. Additionally, they assessed efficiency by incorporating the normalized length of the agent’s path as a weight for the correctness metric.
V-A4 Limitations
@@ V-B Embodied Grasping
Embodied interaction, in addition to question-answering interactions with humans, also involves performing operations based on human instructions, such as grasping and placing objects, thereby completing interactions among the robots, humans and objects. Embodied grasping requires comprehensive semantic understanding, scene perception, decision-making, and robust control planning. The embodied grasping methods integrate traditional robotic kinematic grasping with large models such as LLMs [280] and vision-language foundation models [14], which enables agents to perform grasping tasks under multi-sensory perceptions, including visual active perception, language understanding and reasoning. Figure 12 (b) illustrates an overview of human-agent-object interaction, where the agent accomplishes embodied grasping tasks.
TABLE X: Embodied grasping datasets.
@@
V-B1 Gripper
The current research focus in grasping technology is on two-finger parallel grippers and five-finger dexterous hands. For two-finger parallel grippers, grasping postures are generally categorized into two types: 4-DOF and 6-DOF[290]. The 4-DOF grasp synthesis [286, 282, 281] defines the grasp using a three-dimensional position and a top-down hand orientation (yaw), commonly referred to as “top-down grasping”. In contrast, 6-DOF grasp synthesis [291, 292, 284] defines the grasp posture through a six-dimensional position and orientation. For five-finger dexterous hand grippers, the ShadowHand, a widely used five-finger robotic dexterous hand, features 26 degrees of freedom (DOF). This high dimensionality significantly increases the complexity of generating effective grasp postures and planning execution trajectories.
V-B2 Datasets
Recently, a substantial number of grasping datasets [281, 282, 283, 284, 285] have been generated. These datasets typically contain annotated grasping data based on images (RGB, depth), point clouds, or 3D scenes. With the advent of MLMs and the application of foundational language models to robotic grasping, there is an urgent need for datasets that include linguistic text. Consequently, existing datasets have been extended or reconstructed to create semantic-grasping datasets [287, 288, 293, 289]. These datasets are instrumental in studying grasping models grounded in language, enabling agents to develop a broad understanding of semantics.
Traditional grasping datasets encompass data for both single objects [281] and cluttered scenes [286], providing stable grasp annotations (4-DOF or 6-DOF) that conform to kinematics for each object. These data can be collected from real desktop environments [281], typically including RGB, depth, and point cloud data, or from virtual environments [284], which include image data, point clouds, or scene models. While these datasets are useful for grasping models, they lack semantic information. To bridge this gap, these datasets have been augmented or extended with semantic expressions [287, 294], thereby linking language, vision, and grasping. By incorporating semantic information, agents can better understand and execute grasping tasks. This enhancement allows for the development of more sophisticated and semantically aware grasping models, facilitating more intuitive and effective interaction with the environment. Table X presents the datasets described above, including traditional grasping datasets and language-based grasping datasets.
V-B3 Language-guided grasping
@@
V-B4 End-to-End Approaches
CLIPORT [294] is a language-conditioned imitation learning agent that combines the vision-language pre-trained model CLIP with the Transporter Net to create an end-to-end dual-stream architecture for semantic understanding and grasp generation. It is trained using a large number of expert demonstration data collected from virtual environments, enabling the agent to perform semantically guided grasping. Based on the OCID dataset, CROG [287] proposes a vision-language-grasping dataset and introduces a competitive end-to-end baseline. It leverages CLIP’s visual foundation capabilities to learn grasp synthesis directly from image-text pairs. Reasoning Grasping [288] introduces the first reasoning grasping benchmark dataset based on the GraspNet-1 Billion dataset and proposes an end-to-end reasoning grasping model. The model integrates multimodal LLMs with vision-based robotic grasping frameworks to generate grasps based on semantics and vision. SemGrasp [289] is a method for semantic-based grasp generation that incorporates semantic information into grasp representations to generate dexterous hand grasp postures. It introduces a discrete representation aligning grasp space with semantic space, enabling the generation of grasp postures according to language instructions. To facilitate training, a large-scale grasp-text alignment dataset CapGrasp is proposed.
V-B5 Modular Approaches
@@
These approaches advance the field of language-guided grasping by leveraging both end-to-end and modular frameworks, thereby enhancing the ability of robotic agents to understand and execute complex grasping tasks through natural language instructions. Embodied grasping allows robots to interact with objects, thus improving their intelligence and utility in home services and industrial manufacturing. However, existing embodied grasping methods have limitations, such as reliance on extensive data and poor generalization to unseen data. Future research will focus on improving the generality of agents, enabling robots to understand more complex semantics, grasp a wider variety of unseen objects, and complete intricate grasping tasks.
Figure 13: Chronological overview of the embodied agents. Different colors show different paradigms. MLM means the Multimodal Language Model which directly perceive the world and control the embodiment, VLM means the Visual-Language Model with the outer policy models, LLM + VLM means the LLM-based agent that perceives the world utilizing the VLM, and LLM means the Large-Language Model with visual context and outer policy models.
@@ VI Embodied Agent
An agent is defined as an autonomous entity capable of perceiving its environment and acting to achieve specific objectives. Recent advancements in MLMs have further expanded the application of agents to practical scenarios. When these MLM-based agents are embodied in physical entities, they can effectively transfer their capabilities from virtual space to physical world, thereby becoming embodied agents [298]. Fig. 13 shows the chronological overview of embodied agents.
@@
VI-A Embodied Multimodal Foundation Model
Embodied agents are required to recognize their environment visually, understand instructions audibly, and comprehend their own state to enable complex interactions and operations. This demands a model that integrates multiple sensory modalities and natural language processing capabilities to enhance the agent’s understanding and decision-making by synthesizing diverse data types. Thus, the Embodied Multimodal Foundation Model is emerging. Recently, Google DeepMind found that leveraging foundation models and large, diverse datasets is the optimal strategy. They developed a series of works based on the Robotic Transformer (RT) [11], offering substantial insights for future research on embodied agents.
Significant progress has been made in foundational robotics models, evolving from the initial approach in SayCan [299], which used three separate models for planning, affordance, and low-level policy. Q-Transformer [300] later unified affordance and low-level policy, and PaLM-E [301] integrated planning and affordance. Then, RT-2 [302] achieved a breakthrough by consolidating all three functions into a single model, enabling joint scaling and positive transfer. This represents a substantial advancement in robotics foundational models. RT-2 introduced the Vision-Language-Action (VLA) model, featuring “chain-of-thought” reasoning abilities that enable multi-step semantic reasoning, such as selecting alternative tools or beverages in various contexts. Ultimately, RT-H [4] achieved an end-to-end robot transformer with action hierarchies, to reason about the task planning at a fine-grained level.
To address the generalization limitations of embodied models, Google collaborated with 33 leading academic institutions to create the comprehensive Open X-Embodiment dataset [303], integrating 22 diverse data types. Using this dataset, they trained the universal large model RT-X. This has also promoted the participation of more open-source VLMs for robotics, such as EmbodiedGPT [304] based on LLaVA and RoboFlamingo [305] based on Flamingo. Although Open X-Embodiment provides a vast array of datasets, constructing datasets remains a challenge given the rapid evolution of embodied robotic platforms. To address this issue, AutoRT [306] created a system for deploying robots in new environments to collect training data, leveraging LLMs to enhance learning capabilities through more comprehensive and diverse data.
Additionally, transformer-based architectures face inefficiency problems because embodied models require long contexts that include information from vision, language, and embodied states, as well as memory related to the currently executed tasks. For instance, RT-2, despite its strong performance, has an inference frequency of only 1-3Hz. Several efforts have been made, such as deploying models through quantization and distillation. Moreover, improving the model framework is another viable approach. SARA-RT [307] employs more efficient linear attention, while RoboMamba [308] utilizes the mamba architecture, which is more efficient for long-sequence tasks. This enables it to achieve inference speeds seven times faster than existing robotic MLMs.
Generative-model-based RT excels in high-level task understanding and planning but has limitations in low-level action planning due to the generative model’s inability to precisely generate action parameters and the gap between high-level task planning and low-level action execution. To address this, Google introduced RT-Trajectory [309], which provides low-level visual cues for learning robot control strategies by automatically adding robot trajectories. Similarly, building on the RT-2 framework, the Robot Transformer with Action Hierarchies (RT-H) incorporates a hierarchical action framework, linking high-level task descriptions with low-level robot motions through intermediate linguistic actions [4]. Furthermore, VLA models exhibit emergent capabilities only in high-level planning and affordance tasks related to VLMs. They fail to demonstrate new skills in low-level physical interactions and are constrained by the skill categories in their datasets, resulting in clumsy actions. Future research should integrate reinforcement learning into the training framework of large models to improve generalization, enabling VLA models to autonomously learn and optimize low-level physical interaction strategies in real-world environments, thus executing various physical actions more dexterously and accurately.
VI-B Embodied Task Planning
As previously discussed, a task “put an apple on a plate”, the task planner will divide it into sub-tasks “find the apple, pick the apple”, “find the plate”, “put down the apple”. Since how to find (navigation task) or pick/put down actions (grasping task) are not within the scope of task planning. These actions are typically predefined within simulators or executed in real-world scenarios using pre-trained policy models, such as using CLIPort[294] for grasping tasks.

### v8

@@ V Embodied Interaction
Embodied interaction refer to scenarios where agents interact with humans and the environment in physical or simulated space. The typical embodied interaction tasks are Embodied Question Answering (EQA) and embodied grasping.
@@ V-A1 Datasets
We briefly introduce several embodied question answering datasets, which are summarized in Table VI.
TABLE VI: Comparison of Different EQA Datasets.
@@
EXPRESS-Bench[148] 2025 Exploration-aware EQA HM3D Habitat VLM open-ended 2,044
EQA v1[138] is the first dataset designed for EQA. Built on synthetic 3D indoor scenes from the SUNCG dataset[73] within the House3D[149] simulator, EQA v1 comprises four types of questions: location, color, color_room, and preposition. Similar to EQA v1, MT-EQA[139] is built in House3D using SUNCG by executing functional programs consisting of some basic operations. However, it further extends the single-object question answering task to a multi-object setting. Six types of questions are designed, involving the comparison of color, distance, and size between multiple objects. MP3D-EQA[140] is built on a simulator developed based on MINOS[150] using the Matterport3D dataset[151], expanding the question-answering task to a realistic 3D environment. Referring to EQA v1, MP3D-EQA utilizes three types of templates: location, color, and color_room, generating a total of 1,136 questions in 83 home environments. IQUAD V1[141] is built upon AI2-THOR and consists of three types of questions: existence, counting, and spatial relationships. Unlike other datasets, answering IQUAD V1 questions requires the agent to have a good understanding of affordances and interact with the dynamic environment. VideoNavQA[142] decouples the visual reasoning from the navigation aspect of the EQA problem. In this task, the agent accesses videos corresponding to exploration trajectories with sufficient information to answer questions. SQA3D[143] simplifies protocol (QA only) while still preserving the function of benchmarking embodied scene understanding, enabling more complex, knowledge-intensive questions and a much larger scale of data collection.
Unlike previous datasets that explicitly specify target objects in questions, K-EQA[144] features complex questions with logical clauses and knowledge-related phrases, requiring prior knowledge to answer. OpenEQA[145] is the first open-vocabulary dataset for EQA, supporting both episodic memory and active exploration cases. The episodic memory EQA (EM-EQA) tasks involve an agent developing an understanding of the environment from its episodic memory to answer questions. In active EQA (A-EQA) tasks, the agent answers questions by taking exploratory actions to gather necessary information. Utilizing GPT4-V, HM-EQA[146] is constructed in the Habitat simulator using HM3D. It includes 500 questions across 267 different scenes, which can be roughly categorized into identification, counting, existence, status, and location. S-EQA[147] leverages GPT-4 in VirtualHome for data generation and employs cosine similarity calculations to decide whether to retain the generated data, thereby enhancing dataset diversity. In S-EQA, answering questions requires the assessment of a collection of consensus objects and states to reach an existential “Yes/No” answer. EXPRESS-Bench [148] is the largest exploration-aware EQA dataset that consists of 777 exploration trajectories and 2,044 samples. It also introduces novel evaluation metrics to ensure faithful assessment.
@@ V-A2 Methods
The EQA task was first proposed by Das et al.[138]. In their work, the agent consists of four main modules: vision, language, navigation, and answering. These modules are primarily constructed using traditional neural building blocks: Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs). Some subsequent works[152] retained modules like the question answering module proposed by Das et al.[138] and improved the model. Additionally, Wu et al.[152] proposed integrating the navigation and QA modules into a unified SGD training pipeline for joint training, thereby avoiding employing deep reinforcement learning to simultaneously train the separately trained navigation and question answering modules. From the perspective of task singularity, several works[153] expanded the task to include multiple objectives and multi-agent, respectively, making it necessary for the model to store and integrate the information obtained by the agent’s exploration through methods such as feature extraction and scene reconstruction. Considering the interaction between the agent and the dynamic environment, Gordon et al.[141] introduced the Hierarchical Interactive Memory Network. There is also a limitation in previous works where agents are unable to use external knowledge to answer complex questions and lack knowledge of the explored parts of the scene. To address this, Tan et al.[144] leveraged the neural program synthesis method and the table converted from the knowledge and 3D scene graphs, allowing the action planner to access object-related information. Additionally, an approach based on Monte Carlo Tree Search (MCTS) is used to determine the next location for the agent.
@@
LLMs/VLMs Methods. Majumdar et al.[145] used LLMs and VLMs for episodic memory EQA (EM-EQA) and Active EQA (A-EQA) tasks. For EM-EQA task, they considered Blind LLMs, Socratic LLMs with language descriptions of the episodic memory, Socratic LLMs with descriptions of the constructed scene graph, and VLMs processing multiple scene frames. The A-EQA task extended EM-EQA methods with frontier-based exploration (FBE)[154] for problem-independent environment exploration. Some works[146, 155] employed frontier-based exploration method to identify areas for subsequent exploration and to build semantic maps. They ended the exploration early utilizing conformal prediction or image-text matching to avoid over-exploration. Patel et al.[156] emphasized the question answering aspect of the task. They leveraged multiple LLM-based agents to explore the environment and enable them to independently answer questions with “yes” or “no” answers. These individual responses are utilized to train a Central Answer Model, to aggregate responses and generate robust answers.
V-A3 Limitations
@@ V-B Embodied Grasping
Embodied interaction includes not only question-answering but also performing tasks like grasping and placing objects based on human instructions. Embodied grasping combines traditional kinematic methods [157, 158] with large models such as LLMs and vision-language models, enabling multi-sensory perception and reasoning for task execution. Figure 8 (b) shows an overview of human-agent-object interactions where embodied grasping is performed.
TABLE VII: Embodied grasping datasets.
@@
V-B1 Datasets
Recently, a substantial number of grasping datasets [159, 160, 161, 162, 163] have been generated. These datasets typically contain annotated grasping data based on images (RGB, depth), point clouds, or 3D scenes. With the advent of MLMs and the application of foundational language models to robotic grasping, there is an urgent need for datasets that include linguistic text. Consequently, existing datasets have been extended or reconstructed to create semantic-grasping datasets [165, 166, 167]. These datasets are instrumental in studying grasping models grounded in language, enabling agents to develop a broad understanding of semantics.
Traditional grasping datasets encompass data for both single objects [159] and cluttered scenes [164], providing stable grasp annotations (4-DOF or 6-DOF) that conform to kinematics for each object. These data can be collected from real desktop environments [159], typically including RGB, depth, and point cloud data, or from virtual environments [162], which include image data, point clouds, or scene models. While these datasets are useful for grasping models, they lack semantic information. To bridge this gap, these datasets have been augmented or extended with semantic expressions [165, 168], thereby linking language, vision, and grasping. By incorporating semantic information, agents can better understand and execute grasping tasks. This enhancement allows for the development of more sophisticated and semantically aware grasping models, facilitating more intuitive and effective interaction with the environment. Table VII presents the datasets described above, including traditional grasping datasets and language-based grasping datasets.
V-B2 Language-guided grasping
@@
V-B3 End-to-End Approaches
CLIPORT [168] is a language-conditioned imitation learning agent that combines the vision-language pre-trained model CLIP with the Transporter Net to create an end-to-end dual-stream architecture for semantic understanding and grasp generation. It is trained using a large number of expert demonstration data collected from virtual environments, enabling the agent to perform semantically guided grasping. Based on the OCID dataset, CROG [165] proposes a vision-language-grasping dataset and introduces a competitive end-to-end baseline. It leverages CLIP’s visual foundation capabilities to learn grasp synthesis directly from image-text pairs. Reasoning Grasping [166] introduces the first reasoning grasping benchmark dataset based on the GraspNet-1 Billion dataset and proposes an end-to-end reasoning grasping model. The model integrates multimodal LLMs with vision-based robotic grasping frameworks to generate grasps based on semantics and vision. SemGrasp [167] is a method for semantic-based grasp generation that incorporates semantic information into grasp representations to generate dexterous hand grasp postures. It introduces a discrete representation aligning grasp space with semantic space, enabling the generation of grasp postures according to language instructions.
V-B4 Modular Approaches
@@
These approaches advance language-guided grasping by using end-to-end and modular frameworks, enhancing robotic agents’ ability to perform complex grasping tasks from natural language instructions. Embodied grasping improves robots’ intelligence and utility in-home services and industrial manufacturing. However, current methods face limitations, including reliance on extensive data and poor generalization. Future research aims to enhance agent generality, enabling robots to understand complex semantics, grasp a variety of unseen objects, and tackle intricate tasks.
@@ VI Embodied Agent
An agent is defined as an autonomous entity capable of perceiving its environment and acting to achieve specific objectives. Recent advancements in MLMs have further expanded the application of agents to practical scenarios. When these MLM-based agents are embodied in physical entities, they can effectively transfer their capabilities from virtual space to physical world, thereby becoming embodied agents [172].
@@
VI-A Embodied Task Planning